OpenStudy (anonymous):

Jack has a fair coin. He tosses the coin four times. What is the probability that Jack gets heads, then heads, then tails, and then tails? IT HAS TO BE A FRACTION PLZ HELP< I WILL GIVE MEDAL AND FAN!

OpenStudy (anonymous):

@Bookworm14

OpenStudy (anonymous):

plz help!

OpenStudy (perl):

P( HHTT) = P(H) * P(H) * P(T) * P(T)

OpenStudy (anonymous):

i am a tad confused...

OpenStudy (anonymous):

how would i replace the letters with numbers, there are not any

OpenStudy (anonymous):

besides 4, when he said he would flip the coin 4 times

OpenStudy (perl):

we know the probability of heads is 1/2 , on one flip. probability of tails is 1/2, on one flip

OpenStudy (perl):

P( HHTT) = P(H) * P(H) * P(T) * P(T) = 1/2 * 1/2 * 1/2 * 1/2

OpenStudy (anonymous):

okay, i will tell u what i get in a sec...

OpenStudy (perl):

we can do this because each flip is independent

OpenStudy (anonymous):

i got 1/16

OpenStudy (anonymous):

hmmm... now what?

OpenStudy (anonymous):

is that the answer?

OpenStudy (crashonce):

i believe so

OpenStudy (anonymous):

what do u say @perl ?

OpenStudy (perl):

looks good :)

OpenStudy (anonymous):

thats teh answer?

OpenStudy (anonymous):

the

OpenStudy (anonymous):

wow, thx :O

OpenStudy (anonymous):